OpenStudy (anonymous):

The angle of elevation of a ladder leaning against a wall is 45°. If the distance from the base of the ladder to the wall is 12 feet, how long is the ladder

OpenStudy (ghazi):

cos45= base/length of ladder= 12/l now solve for l

OpenStudy (campbell_st):

draw a picture  identify the ratio and work from there...

OpenStudy (ghazi):

0.707=12/L...therefore L= 12/0.707

OpenStudy (anonymous):

cos"45"=(adjecent/hypotynes)=(12/X);then (12/X)=\[(\sqrt{2}/2)\]then "X"=\[12\sqrt{2}\]